# Learning to Route with Confidence Tokens

### v1

@@ 1 Introduction
Given estimates of confidence, a natural follow-up question is how to utilize such estimates to yield desirable system-level behavior. Often, one may have limited computational resources and only be able to locally fine-tune smaller models. However, one may have API access to a larger model with higher base performance for some cost. In this work, we focus on two downstream tasks that utilize this notion of confidence for practical applications: (1) routing, where queries with low confidence are routed to stronger but more costly LLMs, and (2) rejection, where the model may abstain from making a choice, valuable for preventing unsafe behaviors. rejection-to-answer. Our contributions include:
  1. 1.
    Introducing Self-REF, a lightweight fine-tuning strategy that helps LLMs better learn when they should be confident in their predictions
  2. 2.
    Introducing the confidence-based routing setting, where queries that a small local LLM is uncertain about can be optionally routed to a more powerful LLM (at some cost)
  3. 3.
@@
  4. 4.
    Demonstrating that Self-REF outperforms baselines in the both routing and rejection learning settings on four public datasets.
@@ 2.1 Uncertainty Quantification in LLMs
Recent studies on calibrating LLMs have explored leveraging internal features like logits and attention scores to estimate uncertainty without additional training [56, 35, 48]. Post-processing methods such as temperature scaling [14], prompting LLMs to verbalize confidence scores [46], and instruction learning [55, 40, 32, 34] have also shown improvements in calibration. However, uncertainty scores derived from LLMs may exhibit a weak correlation with actual prediction correctness [22]. Thus, while calibration is important, our work goes beyond it and focuses on the downstream correctness and model confidence scores for the purpose of LLM routing and rejection learning.
@@ 2.2 LLM Routing
In confidence-based query-routing scenarios, one line of research trains additional classifiers to route queries to language models (LLMs) based on specific performance metrics and routing data, forming routing pipelines with mixture-of-experts (MoE) systems to boost performance [12, 41, 44, 24]. Another approach leverages routers to optimize utility-efficiency trade-offs by reconstructing model architectures, aiming to reduce inference costs while maintaining performance [8, 54]. In contrast to existing work that focuses on training query routers based on performance metrics, our approach routes queries based on the model confidence of each LLM’s output from the LLMs themselves. The confidence estimations in our method are predicted in autoregressive decoding, making the model’s confidence more closely aligned with the correctness of the output answers. This dynamic approach allows for more accurate and adaptive routing decisions based on real-time confidence assessments.
@@ 3.2 Self-REF
The Self-REF framework fine-tunes a base model $\mathcal{M}{}$ to use confidence tokens and also learn their embeddings. It involves (i) confidence token annotation, where data are augmented with confidence tokens, and (ii) Self-REF fine-tuning on the augmented data, and (iii) confidence score extraction.
@@
Note that the unconfident and confident tokens are generated end-to-end conditioned on both the prefix query and the answer that the LLM itself generated, in contrast to prior work [41, 12, 44] which typically routes based on the query alone with an extra router. Furthermore, by fine-tuning with the LLM’s own predictions on the training set, Self-REF customizes the notions of uncertainty to the LLM itself, rather than to the inherent uncertainty of a query.
@@
After Self-REF training, one can extract a continuous confidence score $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})\in[0,1]$ by getting the token probability of <CN>, normalized over the sum of the <UN> and <CN> token probabilities: $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})=\frac{P(\texttt{<CN>})}{P(\texttt{<UN>})+P(\texttt{<CN>})}$. This continuous score gives us some control over navigating different tradeoffs in downstream settings.
@@ 3.3 Learning to Route and Reject
In confidence-based routing, answers where the smaller fine-tuned local LLM is uncertain, i.e., $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})<t$ for some chosen threshold $t\in[0,1]$, will be routed to a more powerful LLM $\mathcal{M}_{\text{Large}}(\cdot)$. For answers with $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})\geq t$ (i.e., certain), the answers from the local LLM $\mathcal{M}{}(\cdot)$ will be directly returned to the user.
@@ 4.1 Datasets and Baselines
We compare the Self-REF framework to four state-of-the-art baselines for measuring model confidence, following the settings from existing work. Details about the prompts and hyper-parameters utilized in these baselines are given in Appendix E. We list the baselines as follows:
@@ 4.2 Experimental Settings
In the routing task, we aim to accurately route instances that the smaller LLMs (Llama3-8B-Instruct and Mistral-7B-Instruct) struggle with to the larger LLMs (Llama3-70B-Instruct) for improved performance. All experiments utilize Llama3-70B-Instruct with only its strong in-context learning capabilities during instance routing. The routing decisions are determined by the probabilities of the predicted confidence token <CN> and <UN> generated from Self-REF. To characterize the trade-off between cost and performance, we analyze performance along several possible routing thresholds $t$. In particular, for a set of confidence scores $\mathcal{C}_{\mathcal{M},\mathcal{D}}=\{c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})\}_{i=1}^{N}$, we set $t$ at quantiles, where $t=\mathcal{Q}_{p}\left(\mathcal{C}_{\mathcal{M},\mathcal{D}}\right)$ for $p\in\{0,1,2,\dots,20\}$. Instances are routed to the larger LLM when $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})<t$, so these thresholds correspond to routing approximately 0%, 5%, …, 100% of total queries to the larger LLM.
@@
We select Llama-3-8B-Instruct [47] and Mistral-7B-v0.3-Instruct [23] models as the pre-trained local base LLMs for Self-REF. Llama-3 70B model [47] is the larger, more powerful LLM to which unconfident queries are routed. Parameter-efficient fine-tuning is performed using LoRA adapters with ranks 2, 4, and 8 following the observations from [16]. Under at most eight epochs of training, all other hyperparameters on model training are decided by grid search with the perplexity loss on validation sets. More details of hyper-parameters are in Appendix E.
@@ 5.1 Routing Performance (RQ1)
Confidence-based routing using Self-REF consistently achieves the best accuracy vs. routing rate trade-off for all four datasets (MMLU, OpenbookQA, GSM8K, MedQA) and both local LLMs (Llama3-8B-Instruct model and Mistral-7B-Instruct) (see first row of Figure 2). Using Self-REF on Llama3-8B-Instruct for MMLU, confidence-based routing of 39% of queries can achieve comparable performance to routing to the Llama3-70B-Instruct model alone. Similar observations hold for the OpenbookQA, GSM8K, and MedQA datasets, where routing the Llama3-8B-Instruct model with rates 49%, 65%, and 40%, respectively, can yield comparable performance to Llama3-70B-Instruct alone. For Self-REF on Mistral-7B-Instruct, the minimum routing rates to achieve parity with Llama3-70B-Instruct are 70%, 50%, 75%, and 70% for MMLU, OpenBookQA, GSM8K, and MedQA, respectively (see second row of Figure 2)..
Efficiency Gains
Smaller local models often have lower latency (sec) during inference as well as lower monetary cost. Thus, routing only a fraction of queries to larger, more powerful models can significantly improve the efficiency of the overall system compared to relying entirely on the more powerful models. For each small local LLM (Llama3-8B-Instruct and Mistral-7B-Instruct), we select the best routing rate as the minimum routing rate that achieves comparable performance to Llama3-70B-Instruct (see Figure 2 for the trade-off curves). Without sacrificing performance compared to Llama3-70B-Instruct, we observe that Self-REF achieves as much as a 2.03×2.03$\times$2.03 × latency improvement for Llama3-8B-Instruct on MMLU and a 2.00×2.$0$ latency improvement for Mistral-7B-Instruct on OpenbookQA (Table 1).
@@
Figure 3: ROC curves of Llama3-8B-Instruct on MMLU (a) and OpenbookQA (b); and Mistral-7B-Instruct on MMLU (c) and OpenbookQA (d), for the rejection learning task. The true positive rate (recall) indicates the proportion of samples where “none of the above” is the ground truth that are correctly predicted as such. The false positive rate indicates the proportion of original samples that have been falsely predicted for rejection. Across all settings, Self-REF outperforms other baselines in learning to reject.
@@ 5.3 Calibration with Utility (RQ3)
Here, we analyze the correlation between calibration, or how well predicted probabilities align with actual probabilities, and utility for downstream routing. Following the settings from Błasiok et al. [4], Tian et al. [46], three calibration metrics are used to assess the calibration: (1) expected calibration error, ECE [14]; (2) Brier Score, BS [5]; and (3) the cross-entropy score, CE. Alongside these calibration metrics, we also show the rates of routing required for each technique in order to attain comparable performance to Llama3-70B-Instruct. As shown in Table 2, although Self-REF often achieves the best or second-best calibration scores, better-calibrated methods do not guarantee optimal results for routing. This aligns with observations from prior work [22], which found that well-calibrated confidence scores do not necessarily imply a strong correlation with correctness of the predictions. Overall, Self-REF achieves superior routing performance across all datasets and local LLMs, even though it only outperforms baseline methods under most of the calibration assessments.
@@ 5.4 Ablation Study on Gradient Masking
To understand the impact of gradient masking on downstream fine-tuning performance, we perform an ablation study of Self-REF with and without gradient masking on the OpenbookQA dataset. For both Llama3-8B-Instruct and Mistral-7B-Instruct, we observe that gradient masking improves downstream routing task accuracy over not having gradient masking (Figure 4). This aligns with intuition, since the purpose of gradient masking is to prevent the LLM from learning to make wrong predictions when trained on samples with <UN> tokens.
Figure 4: Ablation studies on gradient masking, which benefits the utility of downstream task.
@@ 6 Discussion
For question (1), we believe that one possible reason behind this phenomenon is that fine-tuning with Self-REF does not degrade the model’s performance and effectively introduces correct knowledge, such as niche astronomy knowledge or complicated moral disputes, that the model does not inherently possess. Additionally, since the error distributions of two independent models are unlikely to overlap perfectly, the ability to route can serve as a form of ensembling of models. In the Appendix, we provide some case studies of questions from MMLU that the powerful Llama3-70B-Instruct model answered incorrectly, but Self-REF with the smaller Llama-8B model answered confidently and correctly. As shown in Appendix Figures 6, the Llama3-70B-Instruct model may lack specific knowledge to judge a question of complicated moral disputes without fine-tuning and injecting the knowledge directly. More case studies are provided in Appendix D.
@@ 7 Conclusion
Our findings suggest that integrating confidence tokens into LLMs is a promising step toward improving their reliability and safety, particularly in real-world applications where errors carry significant costs. Future research will focus on integrating explicit rejection mechanisms into the Self-REF fine-tuning process, rather than relying solely on confidence tokens for rejection learning.
@@ B.2 LLM Routing
LLM routing, where all or part of a query is directed from one LLMs to different LLMs, has been studied in the literature in a few capacities: (1) routing entire queries to different LLMs, (2) retrieval-augmented generation (RAG) systems that route part of the query to an external retrieval system, and (3) speculative decoding which utilizes outputs from models of varying complexity in the decoding step.
(1) In the query-routing scenario, one line of work trains additional classification models to route queries to other LLMs for answers based on designated performance metrics  [12, 20, 21] and routing data [41, 44], which eventually form a routing pipeline with mixture-of-experts (MoE) systems [26, 24] to boost the performance. Another line of work leverages routers to measure utility-efficiency trade-offs, aiming to reconstruct model architectures to reduce inference overhead in terms of cost and execution time while maintaining utility [8, 54]. (2) In RAG systems, routing provides effective and efficient navigation to specific data sources, offering efficient and effective indicators for RAG systems to retrieve. A notable approach is Self-RAG [2], which fine-tunes an LLM retriever using augmented data with predefined routing indicators to optimize retrieval performance. This approach improves RAG systems by enabling more efficient retrieval through dynamic routing decisions of each given sentence in the corpus based on predicted routing indicators generated by the fine-tuned LLM retriever. (3) In speculative decoding, the entire decoding process is carried out through an iterative collaboration between a smaller LLM and a larger LLM. Each final output token sequence can be generated by either a smaller LLM or a larger LLM. Specifically, the token decoding process will route to a larger language model when the smaller model either declines to generate or encounters a failure in token generation [6, 31].

### v2

# Learning to Route LLMs with Confidence Tokens

@@ 1 Introduction
Given estimates of confidence, a natural follow-up question is how to utilize such estimates to yield desirable system-level behavior. Often, one may have limited computational resources and only be able to locally fine-tune smaller models. However, one may have API access to a larger model with higher base performance for some cost. In this work, we focus on two downstream tasks that utilize this notion of confidence for practical applications: (1) routing, where queries with low confidence are routed to stronger but more costly LLMs, and (2) rejection, where the model may abstain from making a choice, valuable for preventing unsafe behaviors. Our contributions include:
  • Introducing Self-REF, a lightweight fine-tuning strategy that helps LLMs better learn when they should be confident in their predictions.
  • Introducing the confidence-based routing setting, where queries that a small local LLM is uncertain about can be optionally routed to a more powerful LLM.
@@
  • Demonstrating that Self-REF outperforms baselines in both the routing and rejection learning settings on four public datasets.
@@ 2.1 Uncertainty Quantification in LLMs
Recent studies in LLM uncertainty have explored leveraging internal features like logits and attention scores to estimate uncertainty without additional training [61, 36, 50]. Post-processing methods such as temperature scaling [14], prompting LLMs to verbalize confidence scores [48], and instruction learning/pretraining [60, 42, 32, 35, 10] have also shown improvements in calibration. In contrast to these methods, Self-REF allows the LLM to learn to estimate its confidence at fine-tuning time. Furthermore, this internal notion of confidence is aligned against correctness rather than the consistency of responses after re-sampling, as is studied in [57, 37, 34]. As noted in [22, 51], uncertainty scores derived from LLMs may exhibit a weak correlation with actual prediction correctness, and the most consistent answers are often not the most correct. Finally, while much of the prior work focuses on calibration, our work goes beyond it and focuses on the correlation between downstream utility of model and confidence scores for LLM routing and rejection learning.
@@ 2.2 LLM Routing
In confidence-based query-routing scenarios, one line of research trains additional classifiers to route queries to language models (LLMs) based on observed performance metrics and routing data, forming routing pipelines with trained multi-agent systems or extra routers to boost performance [13, 43, 46, 24, 59]. Another approach leverages LLM cascades, invoking increasing complex LLMs depending on trained scoring functions which predict whether an answer is correct, and consistency of responses upon resampling [8, 58]. Unlike prior work in LLM cascades, Self-REF trains the estimate of the probability of correctness into the LLM itself instead of an external model. Additionally, the estimates of confidence from Self-REF are aligned against the correctness of the predictions, rather than the consistency. Furthermore, Self-REF computes the confidence in the autoregressive decoding step, allowing for LLM itself to condition on its generated output. This dynamic approach allows for more accurate and adaptive routing decisions based on real-time confidence assessments.
@@ 3.2 Self-REF
The Self-REF framework fine-tunes a base model $\mathcal{M}{}$ to use confidence tokens and also learn their embeddings. The framework involves (i) confidence token annotation, where data are augmented with confidence tokens, and (ii) Self-REF fine-tuning on the augmented data, and (iii) confidence score extraction.
@@
Note that the unconfident and confident tokens are generated end-to-end conditioned on both the prefix query and the answer that the LLM itself generated, in contrast to prior work [43, 13, 46] which typically routes based on the query alone with an extra router. By fine-tuning with the LLM’s own predictions on the training set, Self-REF customizes the notion of uncertainty to the LLM itself, rather relying on the inherent uncertainty of a query.
@@
After training, one can extract a continuous confidence score $c_{\mathcal{M}}(x^{(i)},\\ \widehat{y}^{(i)})\in[0,1]$ by getting the token probability of <CN>, normalized over the sum of the <UN> and <CN> token probabilities: $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})=\frac{P(\texttt{<CN>})}{P(\texttt{<UN>})+P(\texttt{<CN>})}$. This continuous score gives us some control over navigating different tradeoffs in downstream settings.
@@ 3.3 Learning to Route and Reject
In confidence-based routing, answers where the smaller fine-tuned LLM is uncertain, i.e., $c_{\mathcal{M}}(x^{(i)},\\ \widehat{y}^{(i)})<t$ for some chosen threshold $t\in[0,1]$, will be routed to a more powerful LLM $\mathcal{M}_{\text{Large}}(\cdot)$. For answers with $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})\geq t$ (i.e., certain), the answers from the local LLM $\mathcal{M}{}(\cdot)$ will be directly returned to the user.
@@ 4.1 Datasets and Baselines
We compare the Self-REF framework to four state-of-the-art baselines for measuring model confidence, following the settings from existing work. Details about the prompts and hyper-parameters utilized in these baselines are given in Appendix D. We list the baselines as follows:
@@ 4.2 Experimental Settings
In the routing task, we aim to accurately route instances that the smaller LLMs (Llama3-8B-Instruct and Mistral-7B-Instruct) struggle with to the larger LLMs (Llama3-70B-Instruct) for improved performance. All experiments utilize Llama3-70B-Instruct with only its strong in-context learning capabilities during instance routing. The routing decisions are determined by the probabilities of the predicted confidence token <CN> and <UN> generated from Self-REF. To characterize the trade-off between cost and performance, we analyze performance along several possible routing thresholds $t$. In particular, for a set of confidence scores $\mathcal{C}_{\mathcal{M},\mathcal{D}}=\{c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})\}_{i=1}^{N}$, we set $t$ at quantiles, where $t=\mathcal{Q}_{p}\left(\mathcal{C}_{\mathcal{M},\mathcal{D}}\right)$ for $p\in\{0,1,2,\dots,20\}$. Instances are routed to the larger LLM when $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})<t$, so these thresholds correspond to routing approximately 0%, 5%, 10%, …, 100% of total queries to the larger LLM.
@@
We select Llama-3-8B-Instruct [49] and Mistral-7B-v0.3-Instruct [23] models as the pre-trained local base LLMs for Self-REF. Llama-3 70B model [49] is the larger, more powerful LLM to which unconfident queries are routed. Parameter-efficient fine-tuning is performed using LoRA adapters with ranks 2, 4, and 8 following the observations from [16]. Under at most eight epochs of training, all other hyperparameters on model training are decided by grid search with the perplexity loss on validation sets. More details of hyper-parameters are in Appendix D.
@@ 5.1 Routing Performance (RQ1)
Confidence-based routing using Self-REF consistently achieves the best accuracy vs. routing rate trade-off for all four datasets (MMLU, OpenbookQA, GSM8K, MedQA) and both local LLMs (Llama3-8B-Instruct model and Mistral-7B-Instruct) (Figure 2). Using Self-REF on Llama3-8B-Instruct for MMLU, confidence-based routing of just 39% of queries can achieve comparable performance to routing to the Llama3-70B-Instruct model alone. Similar observations hold for the OpenbookQA, GSM8K, and MedQA datasets, where routing the Llama3-8B-Instruct model with rates 49%, 65%, and 40%, respectively, can yield comparable performance to Llama3-70B-Instruct alone. For Self-REF on Mistral-7B-Instruct, the minimum routing rates to achieve parity with Llama3-70B-Instruct are 70%, 50%, 75%, and 70% for MMLU, OpenBookQA, GSM8K, and MedQA datasets, respectively.
@@
Efficiency Gains
Smaller local models often have lower latency (sec) during inference as well as lower monetary cost. Thus, routing only a fraction of queries to larger, more powerful models can significantly improve the efficiency of the overall system compared to relying entirely on the more powerful models. For each small local LLM (Llama3-8B-Instruct and Mistral-7B-Instruct), we select the best routing rate as the minimum routing rate that achieves comparable performance to Llama3-70B-Instruct (see Figure 2 for the trade-off curves). Without sacrificing performance compared to Llama3-70B-Instruct, we observe that Self-REF achieves a 2.03×2.03$\times$2.03 × latency improvement for Llama3-8B-Instruct on MMLU and a 2.00×2.$0$ latency improvement for Mistral-7B-Instruct on OpenbookQA datasets (as shown in Table 1).
@@
Figure 3: ROC curves of Llama3-8B-Instruct on MMLU (a) and OpenbookQA (b); and Mistral-7B-Instruct on MMLU (c) and OpenbookQA (d), for the rejection learning task. The true positive rate (recall) indicates the proportion of samples where “none of the above” is the ground truth that are correctly predicted as such. The false positive rate indicates the proportion of original samples that have been falsely predicted for rejection. Across all settings, Self-REF outperforms other baselines.
@@ 5.3 Calibration with Utility (RQ3)
Here, we analyze the correlation between calibration, or how well predicted probabilities align with actual probabilities, and utility for downstream routing. Following the settings from Tian et al. [48], Błasiok et al. [4], three calibration metrics are used to assess the calibration: (1) expected calibration error, ECE [14]; (2) Brier Score, BS [5]; and (3) the cross-entropy score, CE. Alongside these calibration metrics, we also show the least rates of routing required for each technique to attain comparable performance to Llama3-70B-Instruct. As shown in Table 2, although Self-REF achieves the best or second-best calibration scores, better-calibrated methods do not guarantee optimal results for routing. This aligns with observations from prior work [22], which found that well-calibrated confidence scores do not necessarily imply a strong correlation with predictions correctness. Overall, Self-REF achieves superior routing performance across all datasets and local LLMs, even though it only outperforms baseline methods under most of the calibration assessments.
@@
Figure 4: Ablation studies on gradient masking among Llama3-8B-Instruct and Mistral-7B-Instruct demonstrate its benefits for improving the utility of downstream tasks.
@@ 5.4 Ablation Study on Gradient Masking
To understand the impact of gradient masking on downstream fine-tuning performance, we perform an ablation study of Self-REF with and without gradient masking on the OpenbookQA dataset. For both Llama3-8B-Instruct and Mistral-7B-Instruct, we observe that gradient masking improves downstream routing task accuracy over not having gradient masking (Figure 4). This aligns with intuition since the purpose of gradient masking is to prevent the LLM from learning to make wrong predictions when trained on samples with <UN> tokens.
5.5 Transferability of Confidence Token
We further explore the transferability of Self-REF across unseen datasets (i.e., not included in the training data). With fine-tuned LoRA weights on OpenbookQA, we test on MMLU datasets for Llama3-8B-Instruct and Mistral-7B-Instruct. The results show that Self-REF outperforms the baselines with good routing performance on the MMLU dataset even when using the transferred LoRA weights of the OpenbookQA dataset (Figure 5). To achieve parity with the 70B model, we observe that 75% and 70% of routing rate are required on MMLU datasets under Mistral-7B-Instruct and Llama3-8B-Instruct, respectively. Furthermore, compared to the results obtained by direct fine-tuning on MMLU (i.e., a 70% routing rate on Mistral-7B-Instruct and a 40% routing rate on Llama3-8B-Instruct shown in Figure 2), transferred Self-REF achieves competitive routing performance with only a slight degradation in routing rates. This demonstrates the potential of Self-REF to be trained for general, multi-task purposes due to its transferability.
@@ 6 Discussion
For question (1), we believe that one possible reason behind this phenomenon is that fine-tuning with Self-REF does not degrade the model’s performance and effectively introduces correct knowledge, such as niche astronomy knowledge or complicated moral disputes, that the model does not inherently possess. Additionally, since the error distributions of two independent models are unlikely to overlap perfectly, the ability to route can serve as a form of ensembling of models. In the Appendix, we provide some case studies from MMLU that the powerful Llama3-70B-Instruct model answered incorrectly, but Self-REF with the smaller Llama-8B model answered confidently and correctly. As shown in Appendix Figures 6, the Llama3-70B-Instruct may lack specific knowledge to judge a question of complicated moral disputes without fine-tuning and injecting the knowledge directly. More case studies are provided in Appendix 5.5.
@@
Through uncertainty-based routing or rejection, Self-REF is designed to improve overall system performance while reducing cost. In our analysis, we primarily consider cost through the lens of GPU-hours and latency during inference, ignoring network signal transmission times between different models on different devices. While these could also factor into the cost-performance tradeoffs, signal quality across devices depends on the congestion and bandwidth of the network, and so it is difficult to reliably account for it in the analysis. Another potential risk of routing to larger LLMs is the possibility of data leakage during transmission across the network. Routing user queries to different LLMs can also come at a cost to privacy, especially when data is transmitted to third-party or unauthorized models for inference. Further research into securing routing paths and protocols for communication between devices on private cloud servers could help reduce this type of routing cost.
@@ 7 Conclusion
Our findings suggest that integrating confidence tokens into LLMs is a promising step toward improving their reliability and safety, especially in real-world applications where errors carry significant costs. Future research could study explicit incorporation of a rejection choice into the Self-REF fine-tuning process, rather than relying solely on thresholding confidence token scores for rejection learning.
@@ B.2 LLM Routing
LLM routing, where all or part of a query is directed from one LLMs to different LLMs, has been studied in the literature in a few capacities: (1) routing entire queries to different LLMs, (2) retrieval-augmented generation (RAG) systems that route part of the query to an external retrieval system, and (3) speculative decoding which utilizes outputs from models of varying complexity in the entire decoding step.
(1) In the query-routing scenario, one line of work trains additional classification models to route queries to other LLMs for answers based on designated performance metrics  [13, 20, 21] and routing data [43, 46], which eventually form a routing pipeline with mixture-of-experts (MoE) systems [26, 24] to boost the performance. Another line of work leverages routers to measure utility-efficiency trade-offs, aiming to reconstruct model architectures to reduce inference overhead in terms of cost and execution time while maintaining utility [8, 58]. (2) In RAG systems, routing provides effective and efficient navigation to specific data sources, offering efficient and effective indicators for RAG systems to retrieve. A notable approach is Self-RAG [2], which fine-tunes an LLM retriever using augmented data with predefined routing indicators to optimize retrieval performance. This approach improves RAG systems by enabling more efficient retrieval through dynamic routing decisions of each given sentence in the corpus based on predicted routing indicators generated by the fine-tuned LLM retriever. (3) In speculative decoding, the entire decoding process is carried out through an iterative collaboration between a smaller LLM and a larger LLM. Each final output token sequence can be generated by either a smaller LLM or a larger LLM. Specifically, the token decoding process will route to a larger LLM when the smaller model either declines to generate or encounters a failure in token generation [6, 31].